OpenStudy (goformit100):

Consider the following information regarding the number of men and women workers in three factories I, II and III Men workers W omen workers I30 25 II 25 31 III 27 26 Represent the above information in the form of a 3 × 2 matrix. What does the entry in the third row and second column represent?

OpenStudy (goformit100):

@Jonask

OpenStudy (anonymous):

\[3 \times 2 \implies \text{for rows we put factories and for coulumns gender number} \]

OpenStudy (goformit100):

Thank you sir

OpenStudy (anonymous):

\[\left[\begin{matrix}30 & 25 \\ 25 & 31\\27&26\\\end{matrix}\right]\]

OpenStudy (anonymous):

26 is the number of women in factory three III i think it has numerical improtance